below is a random list of items that belong on either an agribusinesses balance sheet on usinesses balance sheet or income s$4000 “others” is an expense
Answer #1


March 2016 March, 16 và u u 4 01 Tuesday (061-305) Wk 10 6 7 8 9 13 14 15 16 20 21 22 23 27 28 29 30 6 boross profil - sale -28 29 Wednesday 02 (062-304) Wk 10 Earnings Before Interest and Taxes (EBIT) biross proht - Tofal operating - - Interest - 10Tuesday (061-305) Wk 10 27 28 29 30 6 boross profil - sale - cost of goods sold - 15,800,000 - 14,74200 = 1058,000 1 - GrossTIJ (063-303) Wk 10 08 Total current Assets 55+9,000 49,31,000 other current Assets Account Receivable Inventory cash 84, 83,27 28 29 30 Friday 04 (064-202) w 10 Total = Assets Total current Assets t Total Non- 9 ewulent Assets 1,92,08,000 + 2,01, 9921 22 23 28 29 30 Friday 04 Total = Assets Total current Assets Total Non- 9 ewutent Assets 1,92,08,000 + 2,01, 99,000 = 3,94U Saluruy (065-301) Wk 10 - facity (balay ty/balancing 2,03, 60,000 figurer Net Income 3,27,000 Total Non current 89,70,ooo.19 1920 262 26 27 28 29 30 Monday UT (067-299) Wk 11 Income mund 96,000 Earning after 327,000 Tax / Met income 4 Assuming Int